# Chain-of-Action: Trajectory Autoregressive Modeling for Robotic Manipulation

### v1

@@ 3 Chain-of-Action for robotic manipulation
$p(\mathbf{a}_{1:T}\mid\mathbf{O})=p(\mathbf{a}_{T}\mid\mathbf{O})\cdot p(\mathbf{a}_{T-1}\mid\mathbf{O},\mathbf{a}_{T})\cdot p(\mathbf{a}_{T-2}\mid\mathbf{O},\mathbf{a}_{T-1:T})\cdot\cdots\cdot p(\mathbf{a}_{1}\mid\mathbf{O},\mathbf{a}_{2:T})\,,$ (1)
@@
1 Inputs: dataset $\mathcal{D}$
@@ 13      Update $\theta$, $x_{\text{SOS}}$ via backprop on $\mathcal{L}_{\text{total}}$
1 Inputs: image $\mathbf{I}$, proprioceptive state $\mathbf{S}$
@@ 7      Append $\hat{x}_{t}$ to $H$
Locality modeling  Multi token prediction (MTP) [7] can serve as a regularization for action locality modeling. We assign the last $K$ layers of the transformer decoder to produce predictions for different future steps. Concretely, layer $k$ predicts token $\hat{x}_{t+k}$, where $k=1,...,K$, making the model aware of the mutual dependencies across the next $K$ steps in a single forward pass. This design introduces temporal locality into the decoding process, enhancing stability in long-horizon generation while remain our global-to-local chain-like structure. Importantly, this regularization is applied only during training and removed at inference.
@@ 4 Implement details
where $\hat{\mathbf{a}}_{t+k-1}^{k}$ and $\hat{x}_{t+k-1}^{k}$ are the predicted action and its latent embedding from $k$-th head of MTP layer at step $t$, and $f_{\text{enc}}(\cdot)$ is the action encoder network. Note that for decoding steps where $t+k-1>T$, the corresponding terms are masked out and do not contribute to the loss. This ensures that predictions beyond the trajectory horizon are excluded from supervision. For parallel training with a batch of samples, we set $T_{\text{max}}$ as the maximum sub-trajectory length (practically the longest in the dataset), and zero-pad all shorter sequences accordingly. The loss for padded steps is masked out to avoid affecting gradient updates.
@@ 5.1 Simulation experiment settings
Baseline  We compare our method against three representative visuo-motor policies: ACT, Diffusion Policy (DP), and Octo [28]. For ACT and DP, we follow the RLBench training protocol introduced in [33]. To better show the effectiveness of our modeling paradigm, we align our basic architecture with ACT, applying modifications mainly in the transformer decoder, as detailed in Section 4. For Octo, we adopt the official fine-tuning instructions. Note that Octo benefits from large-scale pretraining, and is therefore not directly comparable to our method in terms of data scale. The performance of these baselines, for example perfect results on tasks such as Sweep Dust and competitive performance on others (see Table 1), confirms that all reference models are properly trained.
Training and evaluation protocol  Each method is trained on 100 demonstrations per task. Although RLBench tasks include multiple variations, we restrict training to variation 0 to reduce computational cost. Evaluation is mainly conducted on 60 tasks compared with ACT and DP, each with 25 demonstrations. A subset of 10 widely used tasks (refers to Table 1) is used for comparison with Octo and ablation study.
@@
Figure 4: Success rate improvement across 60 RLBench tasks, sorted by improvement from high to low. The average success rate over all tasks is shown in the inset on the right.
@@ 5.2 Overall comparisons
The overall results are presented in Figure 4, with task-wise averages summarized in the accompanying wrapped table. To better assess the effectiveness of our method, we report task-level improvements over both ACT and DP. Compared to ACT, our method achieves higher success rates on 81.7% of the tasks, with an average improvement of 16.3%. Relative to DP, our method improves performance on 80.0% of the tasks, with an average gain of 23.2%. These improvements are especially pronounced in tasks involving significant variation in object position and pose, indicating stronger spatial generalization. As ACT and CoA share a consistent Transformer encoder-decoder architecture and are trained on the same setting, the observed gains highlight the effectiveness of our modeling paradigm. The results suggest that a principled change in how action sequences are represented and generated can lead to substantially better performance under distribution shifts.
In addition, the complementary comparison with Octo and detailed results with ACT and DP over 10 selected tasks are shown in Table 1. Visualization of generated trajectories are shown in Figure 3.
@@ 5.3 Dive into spatial generalization
Although CoA significantly outperforms ACT and DP on the overall benchmark, it remains crucial to understand why such improvements emerge. To this end, we investigate the spatial generalization behavior of our model from three complementary perspectives:

### v2

@@ 3 Chain-of-Action for robotic manipulation
$p(a_{1:T}\mid O)=\underbrace{p(a_{T}\mid O)}_{\text{Keyframe Action}}\cdot\underbrace{p(a_{T-1}\mid a_{T},O)\dots p(a_{2}\mid a_{3:T},O)}_{\text{Reverse Reasoning Actions}}\cdot\underbrace{p(a_{1}\mid a_{2:T},O)}_{\text{Executed Action}}$ (1)
@@
1Inputs: dataset $\mathcal{D}$
@@ 13 Update $\theta$, $x_{\text{SOS}}$ via backprop on $\mathcal{L}_{\text{total}}$
1Inputs: image $\mathbf{I}$, proprioceptive state $\mathbf{S}$
@@ 7 Append $\hat{x}_{t}$ to $H$
Locality modeling  Multi token prediction (MTP) [7] can serve as a regularization for action locality modeling. We assign the last $K$ layers of the transformer decoder to produce predictions for different future steps. Concretely, layer $k$ predicts token $\hat{x}_{t+k}$, where k=1,,Kk=1,...,K, making the model aware of the mutual dependencies across the next $K$ steps in a single forward pass. This design introduces temporal locality into the decoding process, enhancing stability in long-horizon generation while remain our global-to-local chain-like structure. Importantly, this regularization is applied only during training and removed at inference.
@@ 4 Implement details
where $\hat{\mathbf{a}}_{t+k-1}^{k}$ and $\hat{x}_{t+k-1}^{k}$ are the predicted action and its latent embedding from $k$-th head of MTP layer at step $t$, and $f_{\text{enc}}(\cdot)$ is the action encoder network. Note that for decoding steps where t+k1>Tt+k-1>T, the corresponding terms are masked out and do not contribute to the loss. This ensures that predictions beyond the trajectory horizon are excluded from supervision. For parallel training with a batch of samples, we set $T_{\text{max}}$ as the maximum sub-trajectory length (practically the longest in the dataset), and zero-pad all shorter sequences accordingly. The loss for padded steps is masked out to avoid affecting gradient updates.
@@ 5.1 Simulation experiment settings
Baseline  We compare our method against representative approaches from three categories: (1) training visuomotor policies from scratch, including ACT and Diffusion Policy (DP); (2) finetuned generalist robotic policies, represented by Octo [29]; (3) 3D-based hierarchical methods, including PerAct [35], 3D Diffuser Actor [17], and RVT-2 [9]. We note 3D-based hierarchical methods fundamentally differ from our approach by relying on 3D inputs and motion planners to generate trajectories. We provide additional discussion on these differences in Appendix 7.
Training and evaluation protocol  To ensure broad and representative evaluation, our main benchmark is conducted on a tailored set of 60 RLBench tasks, where CoA is compare with ACT and DP, Each method is trained on 100 demonstrations and evluated on 25 demonstrations per task. For ACT and DP, we follow the RLBench training protocol introduced in [34], which is detailed in Appendix 10. To better demonstrate the effectiveness of our modeling paradigm, we align our base architecture with ACT and introduce modifications primarily in the transformer decoder, as detailed in Section 4. The strong performance of these baselines—such as perfect success rates on tasks like Sweep Dust and competitive results on others (Table 1)—confirms that all reference models are properly trained. For comparison with Octo, we adopt the evaluation subset RLBench-10 proposed in [44] and use the reported results. This subset is also used for our ablation studies for convenience. To facilitate comparison with 3D-based hierarchical methods, we evaluate on the RLBench-18 split [35], using reported results from prior work [9].
@@
Figure 4: Success rate improvement on RLBench-60, sorted by improvement from high to low. The average success rate over all tasks is shown in the inset on the right.
Table 2: Comparison on the RLBench-18. 3D-based hierarchical methods use 3D point clouds and motion planners, while image-based visuomotor policies operate directly on RGB inputs.
3D-based hierarchical methods Image-based visuomotor policies
Task PerAct 3D Diffuser Actor RVT-2 Image-BC (CNN) Image-BC (ViT) DP ACT CoA
Close Jar 55.2 $\pm$ 4.7 96.0 $\pm$ 2.5 100.0 $\pm$ 0.0 0 0 0 0 0
Drag Stick 89.6 $\pm$ 4.1 100.0 $\pm$ 0.0 99.0 $\pm$ 1.7 0 0 0 0 0
Insert Peg 5.6 $\pm$ 4.1 65.6 $\pm$ 4.1 40.0 $\pm$ 0.0 0 0 0 0 0
Meat off Grill 70.4 $\pm$ 2.0 96.8 $\pm$ 1.6 99.0 $\pm$ 1.7 0 0 16 32 88
Open Drawer 88.0 $\pm$ 5.7 89.6 $\pm$ 4.1 74.0 $\pm$ 11.8 4 0 44 52 88
Place Cups 2.4 $\pm$ 3.2 24.0 $\pm$ 7.6 38.0 $\pm$ 4.5 0 0 0 0 0
Place Wine 44.8 $\pm$ 7.8 93.6 $\pm$ 4.8 95.0 $\pm$ 3.3 0 0 56 56 80
Push Buttons 92.8 $\pm$ 3.0 98.4 $\pm$ 2.0 100.0 $\pm$ 0.0 0 0 0 32 28
Put in Cupboard 28.0 $\pm$ 4.4 85.6 $\pm$ 4.1 66.0 $\pm$ 4.5 0 0 0 0 8
Put in Drawer 51.2 $\pm$ 4.7 96.0 $\pm$ 3.6 96.0 $\pm$ 0.0 8 0 40 60 88
Put in Safe 84.0 $\pm$ 3.6 97.6 $\pm$ 2.0 96.0 $\pm$ 2.8 4 0 24 36 80
Screw Bulb 17.6 $\pm$ 2.0 82.4 $\pm$ 2.0 88.0 $\pm$ 4.9 0 0 0 0 0
Slide Block 74.0 $\pm$ 13.0 97.6 $\pm$ 3.2 92.0 $\pm$ 2.8 0 0 0 36 64
Sort Shape 16.8 $\pm$ 4.7 44.0 $\pm$ 4.4 35.0 $\pm$ 7.1 0 0 0 0 0
Stack Blocks 26.4 $\pm$ 3.2 68.3 $\pm$ 3.3 80.0 $\pm$ 2.8 0 0 0 0 0
Stack Cups 2.4 $\pm$ 2.0 47.2 $\pm$ 8.5 69.0 $\pm$ 5.9 0 0 0 0 0
Sweep to Dustpan 52.0 $\pm$ 0.0 84.0 $\pm$ 4.4 100.0 $\pm$ 0.0 0 0 100 100 92
Turn Tap 88.0 $\pm$ 4.4 99.2 $\pm$ 1.6 99.0 $\pm$ 1.7 8 16 32 36 56
Average 48.7 81.3 81.4 1.33 0.89 17.33 24.44 37.33
@@ 5.2 Overall comparisons
The overall results are presented in Figure 4, with task-wise averages summarized in the accompanying wrapped table. To better assess the effectiveness of our method, we report task-level improvements over both ACT and DP. Compared to ACT, our method achieves higher success rates on 81.7% of the tasks, with an average improvement of 16.3%. Relative to DP, our method improves performance on 80.0% of the tasks, with an average gain of 23.2%. These improvements are especially pronounced in tasks involving significant variation in object position and pose, indicating stronger spatial generalization. As ACT and CoA share a consistent Transformer encoder-decoder architecture and are trained on the same setting, the observed gains highlight the effectiveness of our modeling paradigm. The results suggest that a principled change in how action sequences are represented and generated can lead to substantially better performance under distribution shifts. The detailed per-task results are in Appendix 8. In addition, the comparison with Octo and detailed results with ACT and DP over 10 selected tasks are shown in Table 1. Results on the RLBench-18 benchmark are reported in Tab 2. We observe that, in these settings, CoA outperforms the finetuned generalist robot policy Octo, while a substantial performance gap remains compared to the 3D-based hierarchical methods.
@@ 5.3 Dive into spatial generalization
Although CoA significantly outperforms ACT and DP on the overall benchmark, it remains crucial to understand why such improvements emerge. To this end, we investigate the spatial generalization behavior of our model from three complementary perspectives.
First, in the Interpolation vs. Extrapolation case study, we analyze CoA’s performance under controlled spatial distributions within a single representative task. This study reveals that CoA not only achieves higher success rates under in-distribution (interpolation) configurations, but also demonstrates a substantially larger advantage in out-of-distribution (extrapolation) settings, indicating stronger spatial generalization.
Second, in Correlation with spatial distribution, we quantitatively examine how task performance correlates with spatial variation difficulty across the 60 RLBench tasks. The results show that CoA consistently improves over ACT and DP across all spatial variance levels, and that the performance gap widens as spatial generalization becomes more challenging.
Finally, in Attention-based analysis of action chain, we visualize the attention maps between action tokens in the Transformer decoder. The attention patterns clearly reveal structured dependencies along the predicted action sequence, supporting the hypothesis that CoA performs chain-like global-to-local reasoning throughout the trajectory generation process.